OpenStudy (anonymous):

if anyone could please answer these 2 problems ill give you a medal .. please answer asap ! Shawn and Ellie are construction workers. Shawn can complete a concrete job in 4 hours, while Ellie can complete it in 6 hours. The foreman says that it will take them 5 hours to complete it if they work together. Explain each step in solving this equation, and determine if the foreman is correct or not. (10 points) Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences,explain

OpenStudy (anonymous):

is this multiple choice or not

OpenStudy (anonymous):

If Shawn works for one hour he will complete 1/4 of the job. If Ellie works for one hour she will complete 1/6 of the job. If they work together for an hour they will complete 1/4 + 1/6 of the job. 14+16=512 ..........(1) So, working together for one hour, they will complete 5/12 of the job. Let x be the number of hours working together need to complete the job. Then the following equation can be written: 5x12=1 ...............(2) Now you just need to solve equation (2) to find the time needed to complete the job when the two people work together.

OpenStudy (anonymous):

The answer is 125=2.4 hours or 2 hours 24 minutes

OpenStudy (anonymous):

are you still there @Gerial

OpenStudy (anonymous):

yes i am they are not multiple choice you have to explain how you got your answer an work @ogjb_

OpenStudy (anonymous):

yh i sent it in

OpenStudy (anonymous):

do you need help with the second

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

yes i dont know what to do at all but i will give you your medal now

OpenStudy (anonymous):

@ogjb_

OpenStudy (anonymous):

ill help you with second

OpenStudy (anonymous):

ok i have a question when you typed the equations then put "5x12=1 ...............(2) " is that part of the answer or ?

OpenStudy (anonymous):

ok let me see backward that should like x(x-8)/(x+x-8) = 3 with x is the Jimmy's time if he complete the job alone x(x-8)/(x+x-8) = 3 (x^2 - 8x)/(2x-8) = 3 cross multiply then simplify and solve for x (notice that it must be positive) x2−8x2x−8=31 x^2-8x = 6x-24 simply by combine then solve for x x^2-14x+24=0? (x-12)(x-2)=0 x=12 and x=2 but we have to recheck again which one of that numbers can be solution for this problem. see before i take by assumed x = jimmy's time and x-8 = harry's time if they do the job alone. now if you subtitute both value of x above, which one possible as the answer ? if x = 2, then x - 8 = 2 - 8 = -6 (it cant) if x = 12, then x-8 = 12-8 = 4 (nah, this is a positive number), so x = 12 is satisfies as the solution so x=12 is the solution it would give us Jimmy = 12 Harry = 4 (see that harry 8 hours faster than Jimmy

OpenStudy (anonymous):

did you see the work

OpenStudy (anonymous):

cross multiply then simplify and solve for x (notice that it must be positive) x2���������������������������������������������������������������������������������8x2x���������������������������������������������������������������������������������8=31

OpenStudy (anonymous):

@ogjb_ what is that & how do i give you an medal ?

OpenStudy (anonymous):

hold on @Gerial let me see if i can send it again im sorry

OpenStudy (anonymous):

ok i have i now ! thank you so much ..

OpenStudy (anonymous):

x2−8x2x−8=31 so x^2-8x/6x-24? x^2-8x = 6x-24 simply by combine then solve for x x^2-14x+24=0? (x-12)(x-2)=0 then set up each factor be equal zero then solve for x x=12 and x=2 ok so x=12 is the solution because it would give us Jimmy = 12 Harry = 4 (see that harry 8 hours faster than Jimmy)

OpenStudy (anonymous):

idk why it keeps doing that

OpenStudy (anonymous):

x2−8x2x−8=31

OpenStudy (anonymous):

ok this is my last one actually , think you could do it really quick ? @ogjb_ 2.When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = –2, x = 3, and x = 5. Bobby says that the function is undefined at those x